# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

### v1

@@ 8. Experiment & Evaluation
We first describe the experimental setup for model training/testing, and then qualitatively and quantitatively evaluate our method.
@@ 8.1. Experiment Setup
We implement our interactive character control system in Unity 3D engine and update the motion with 60 fps on a computer equipped with 5.1GHz Intel i5-13600KF and NVidia GeForce RTX 4090. For training the neural implicit trajectory planner, we reduce the network size from (Park et al., 2019) to a 6-layer MLP with 256 neurons per layer and a skip connection to the 4-th layer. We prepare 5546 interactions, each of which is composed of a pick-and-place motion and the local environment geometry represented by position/field-feature pairs § 4.2 uniformly sampled in a resolution of 2.5cm within a $(80\text{cm})^{3}$ cubic region around the target object.
Training the auto-decoder using 90% of our data takes 300 epochs by 8 hours on two Tesla A100s. At runtime, each trajectory planning will optimize the latent code by 100 steps in about 1.25 seconds.
@@
For motion data processing, we extract 8 channels in the frequency domain with the same network structure as (Starke et al., 2022) to train the Phase Auto-encoder. Taking 120 epochs in 10 hours in all the motion data, we then pre-process all the motion data with DeepPhase features. To achieve a real-time inference over 60 fps, we designed the DeepPhase Interaction Controller network with 3 layers of 128 neurons in the Gating Network and the main network with 3 layers of 512 neurons. We train the controller network using 70% of the dataset with 150 epochs, which takes approximately 8 hours.
@@ 8.2. Trajectory Planner Evaluation
Figure 9. Comparison of trajectory planners under representative generalization scenes with narrowed cabinets and expanded obstacles. The blue and green curves shown in two views records the approaching and leaving trajectory of the wrist, respectively. See more comparisons in our supplementary video.
@@ 8.2.1. Comparison on Motion Realism
Comparison to cuRobo: The cuRobo planner aims to produce collision-free motion with an optimal jerk cost; however, its wrist trajectory deviates from realistic human motion in two aspects. First, cuRobo tends to produce straight trajectories, where the hand and the in-hand object move too close to obstacles (see Fig. 9(b,c-2)). Also, as cuRobo is designed for rigid-body grippers, the object tends to produce minimal rotation after grasped, again resulting in little space between the object and the environment. In contrast, our planner, learned from kinematic features, employs dexterous strategies to produce human-like trajectories; for instance, when grasping a spatula out from a drawer, it makes use of the finger degrees of freedom for rotating the in-hand object while pulling it out vertically, thereby reducing the swept volume (see Fig. 9 (a-1)). Secondly, cuRobo can only produce hand trajectories with a fixed base, resulting in a lack of coordination of the locomotion with the pick-and-place task; including the base translation and orientation would be difficult with the minimum jerk criteria as human models require stepping causing many local minima. Indeed, producing human-like motion with optimization in canonical space requires a significant amount of cost/reward tuning in high dimensions, which is not practical. In contrast, our data-driven model motion can produce well-coordinated motions that are well suited for the condition represented by the latent code $\mathbf{z}$ and the condition parameters $\mathbf{c}$, eliminating the fixed hand base to achieve body-level guidance during its trajectory prediction.
Comparison to Diffusion Policy: Although learned from the same kinematic dataset, we find that the Diffusion Policy struggles to robustly produce realistic trajectories in test scenes due to the sparse dataset with high diversity. It suffers from over-straight trajectories (see Fig. 9(a,b-3) blue curves), irregular wrist orientation (see Fig. 9(b,c-3) and sudden change in height resulting in heavy placement (see  Fig. 9(d-3))
Quantitative Evaluation: The quantitative results in Table 4 also validate that our implicit neural trajectory planner can produce body motion more similar to human movement, resulting in an overall 3% lower Fréchet distance to the dataset compared with the motion following the cuRobo trajectory. Moreover, the improved motion smoothness and reduced foot sliding illustrate that our generated trajectory can be better tracked by the character, enabling it to achieve interactions with higher body coordination, benefiting from the aforementioned strengths of our kinematic data-based planner.
Figure 10. Figures from (a) to (d) sequentially show the planning of approaching a bottle with field deformation under a standard-sized shelf, 80% narrow (29.5cm in width) shelf, deep placed in the narrow shelf, and deep placed with another bottle in a narrow shelf.
Figure 11. Visualization of the implicit neural representation of the layouts and approaching trajectories uniformly sampled from the training dataset of Sec. 4.2. We project the latent vector $\mathbf{z}$ of each 3-channel field $\mathbf{D}$ to 2D by t-SNE discarding the condition labels $\mathbf{c}$. The latent space automatically clusters different types of furniture, while similar layouts are continuously distributed. The projection of all the unseen layouts (cases in grey plots) are mapped to those of similar scenes (in orange).
@@ 8.2.2. Evaluation Under Scene Generalization
Although our planner was only trained on standard-sized furniture in our motion capture scenes, it learns the spatial relationships between the trajectories and the environment, and these relationships remain consistent in novel scenes. In Fig.10, we show snapshots of the pick-and-place motions on shelves, along with the adapted trajectories in blue and the predicted time-of-arrival field $\bm{D}_{\text{toa}}$ depicted in green. Under a standard-sized shelf with an ordinary placement (Fig.10(a)), there is ample space for the right hand to reach the goal from the right side. When the shelf’s size is narrowed to 80% of the training size (see Fig.10(b)), the character instead reaches from the front and passes through the small space on the right to grasp the bottle. Further, after setting the bottle’s position deeper into the narrow shelf (Fig.10(c)), the character moves its wrist further to the left before reaching for the object. Upon adding another bottle (Fig. 10(d)), the system further adjusts the hand trajectory to avoid any collision. These experiments demonstrate that our designed three-channel fields of similar layouts can be well-mapped to a continuous latent space by the auto-decoder, enabling our planner to generalize to novel scenes.
@@
To further illustrate the generalization capability, we analyze the latent space of our auto-decoder. Similar to how DeepSDF exhibits continuous latent representations for diverse shapes, our auto-decoder maps diverse scenes and trajectory pairs to the latent space. We project the latent codes onto a plane using t-SNE (Cieslak et al., 2020) and visualize them in Fig. 11. The projected scenes are well clustered, with unseen test scenes mapped adjacent to similar ones. This distribution reveals that even when training on scenes with various heterogeneous furniture together in one model, the auto-decoder structure effectively forms a continuous and smooth latent space for all types of scenes. This allows our runtime optimization to stably find the code that best suits the test environment and conditions, mapping it to a 6D trajectory within the distribution of plausible manipulations.
@@
To quantitatively evaluate the generalization capability of different trajectory planners, we prepare an extended test dataset composed of our test data and those from indoor scene databases  [^2] https://assetstor.unity.com/packages/3d/props/interior/kitchen-props-pack-pbr-204679; featured with various obstacle shapes, layout difference, and placement variations. Similar to Fig. 10, we synthesize unseen layouts by randomly shifting each obstacle object on/inside the involved furniture within 5cm, rotating them within ±$\pm 90^{\circ}$, and re-scaling within 0.81.25×0.8$\sim$ 1.25$\times$0.8 ∼ 1.25 ×. Also, we reshape the containers to new widths, depths, and heights ranging from 70% to 150% of their original dimensions, and remain the pick-able cases.
We compare the trajectory un-smoothness (measured by the mean acceleration), spare safety distance to obstacles from the hand and object, and the success rate among the results from our planner, the NVIDIA cuRobo planner, and the Diffusion Policy, as the examples presented in Tab. 3. Our evaluation benchmark includes 168 densely packed layouts, which will be released alongside our Unity3D project.
@@
As shown in Table 3 part 2, our SE(3) trajectory planner achieves a high success rate over 96%, with a 4.5% higher safety distance than the cuRobo planner and 28.4% higher than the Diffusion Policy, along with significantly better motion smoothness. These results validate the benefits of our planner’s strong generalization capability and more human-like pick-and-place trajectories. The success rate of cuRobo also decreases when the scene is highly cluttered and requires more adjustments to navigate through narrow regions. In these cases, the precise geometric approximation in cuRobo cannot converge, and using a coarser geometric approximation eliminates sufficient constraints, leading to collision-prone policies.
@@ 8.2.3. Ablation Study on Distance Field Channels
We evaluate the effectiveness of our three-field setting by comparing it with a simplified two-channel auto-decoder. In these alternative models, we only keep the goal-reaching distance $D_{t}$ (as INTP-2Ch-$D_{t}$) or we merge the $D_{t}$ together with the obstacle distance $D_{o}$ into one channel $D_{o^{\prime}}$ (as INTP-2Ch-$D_{o^{\prime}}$) by the unsigned/signed distance to the nearest obstacle/target object surface, respectively. As shown in Table 3, the performance decreases significantly when using only two channels. This result confirms that having separate and complete channels for $D_{t}$ and $D_{o}$ is beneficial for high-quality trajectory predictions. Specifically, the $D_{t}$ field away from the object helps to direct the hand to approach, while $D_{o}$ field contributes to collision avoidance.
@@
Table 4. Ablation study and comparison to the human motion data. The unsmoothness is computed by averaging the accelerations among all joints in $cm/s^{2}$. The sliding sums up the feet’ acceleration during ground contact by $cm/s^{2}$. The rooted mean-square curvature (RMSC) of the wrist trajectory is averaged among all the approaching/leaving cases in $cm^{-1}$. The Fréchet distance evaluates the motion similarity to the dataset by extracting each 2s sliding time-windows pose as the feature vector.
[…]
w/ KF + cuRobo 6.54 7.22 691 0.0725
w/o goal phase 6.78 7.55 819 0.0982
w/ matched goal 6.36 6.80 689 0.0721
w/ predicted goal 6.30 6.86 711 0.0614
w/ KF goal 5.94 6.17 670 0.0509
Dataset 2.22 1.02 - 0.0408
Figure 12. Ablation study for our Deep Phase Interaction Controller in typical full-body interactions. We visualize the hands trace for the target approaching at (a,b), where the trajectory pieces with high curvature are highlighted by red heatmaps. In line (c) we show the bipedal trace of a pivot turning case, and line (d) shows a side-stepping case, where the skeletons from green to blue are from sequentially generated key-frames where both the feet are on the ground, with red curves highlighting the sliding pieces. We annotate each approaching trajectory with its RMS curvature in $cm^{-1}$, the mean sliding during stepping in $cm/s$, and the duration time in seconds.
@@ 8.3. Evaluation of the Pick-and-Place Motions
8.3.1. Ablation Study of the Deep Phase Controller
We evaluate the effectiveness of our goal-driven character controller by examining 3 variants of our DeepPhase controller, including a model that drives the character solely based on the goal key joints (w/o goal phase), a model that uses matched goal phase features obtained from the dataset (w/matched goal phase), and a model where the network is fed with goal phase features auto-regressively predicted by itself (w/ predicted goal phase). In addition to testing in the general scene described earlier (see Footnote 2), we use another test scene from our dataset composed of 20 lightweight objects and 9 pieces of furniture. A long motion sequence over 3 minutes is constructed by the user through sequential clicks on objects and furniture. We compute the smoothness, foot skating artifacts, and the Fréchet distance to the corresponding motion in the dataset (see Table 4). The results show that the use of the goal phase improves the accuracy of all measures.
@@ 8.3.1. Ablation Study of the Deep Phase Controller
When no goal phase feature is given, the legs fail to produce effective stepping to let the key joints track their goals, resulting in larger foot skating (see Tab. 4), This is evident in complex locomotion like pivot turning (see Fig.12(c-4)). Also, the upper body can struggle to recover from challenging previous poses without a goal cue, as shown in Fig.12(d-4), where the uncoordinated body causes increased foot sliding.
When the matched goal phase is used instead of that predicted by the network, the system suffers from the poor guidance to the goal. As matched goal phase are those of sparse keyframes, it cannot guide the character to conduct efficient stepping for reaching a state that is spatio-temporally apart (see Fig.12(c-3)). Also, the deviation of the current state and the matched state leads to biased directional guidance, requiring the character to take extra steps/time for reaching the goal (see Fig.12(d-3)).
@@
Furthermore, the trajectory of the end-effector exhibits higher motion realism and smoothness accompanied by body coordination when using the filtered goal phase. Without the goal cue in the frequency domain, tracking only the three key joint goals spatially leads the model to focus excessively on tracking the hip or other hand goals when the manipulation takes place away from the torso. This causes large deviations when approaching the hand goals (see Fig. 12(b-4)) and results in an unsmooth trajectory of the manipulating hand.
Even with matched goal phase priors, the system can produce unsmooth wrist motions when the matched prior suddenly switches from left-hand grasping to right-hand grasping—for example, when the main hand starts to reach out for the target object after the auxiliary hand finishes opening the cabinet, as shown in Fig.12(b-3), and significant direction adjustments occur (see Fig.12(a-3)).
While the approaching motion toward the goal phase directly estimated by the MoE network slightly improves curvature results (see Fig. 12(a-2, b-2)), it still suffers from discrete switches of the goal phase. When moving the hand goal spatially along the planned approaching trajectory, the predicted goal phase updates rapidly to catch up with the goal. In case (b-2), the arm immediately stretches out to reach the target object even before the character is close enough, producing an awkward motion due to overstretching.

### v2

@@ 8. Experiment & Evaluation
We first describe the experimental setup for model training/testing, and then qualitatively and quantitatively evaluate our methods.
@@ 8.1. Experiment Setup
We implement our interactive character control system in Unity 3D engine and update the motion with 60 fps on a computer equipped with 5.1GHz Intel i5-13600KF and NVidia GeForce RTX 4090. For training the neural implicit trajectory planner, we reduce the network size from (Park et al., 2019) to a 6-layer MLP with 256 neurons per layer and a skip connection to the 4th layer. We prepare 5546 interaction cases, each of which is composed of a pick-and-place motion and the local environment geometry represented by position-field feature pairs ( § 4.2) uniformly sampled in a resolution of 2.5cm within a $(80\text{cm})^{3}$ cubic region around the target object.
Training the auto-decoder using 90% of our data takes 8 hours on two Tesla A100s for 300 epochs. At runtime, each trajectory planning task optimizes the latent code by 100 steps on our desktop, taking about 1.25 seconds.
@@
For motion data processing, we extract 8 channels in the frequency domain with the same network structure as (Starke et al., 2022) to train the Phase Auto-encoder. After training it with 120 epochs for 10 hours on the training split, we pre-process the full motion dataset with DeepPhase features. To achieve a real-time inference over 60 fps, we design the DeepPhase interaction control network with 3 layers of 128 neurons in the gating network and the main motion prediction network with 3 layers of 512 neurons. We train the controller network using 70% of the dataset with 150 epochs, which takes approximately 8 hours.
@@ 8.2. Trajectory Planner Evaluation
Figure 9. Comparison of trajectory planners under representative generalization scenes with narrowed cabinets and expanded obstacles. The blue and green curves shown in two views record the approaching and leaving trajectory of the wrist, respectively. See more comparisons in our supplementary video.
@@ 8.2.1. Comparison on Motion Realism
We compare our trajectory planner with two state-of-the-art trajectory planners (see Fig. 9): the parallelized collision-free minimum-jerk model proposed in cuRobo (Sundaralingam et al., 2023), for robotics applications; and the Diffusion Policy (Chi et al., 2023), a generative data-driven planner trained with the same set of motion data by 6000 epochs, taking $\sim$2 days on 2 Tesla V100 GPUs.
Comparison to cuRobo
The cuRobo planner aims to produce collision-free motion with an optimal jerk cost; however, its wrist trajectory deviates from realistic human motion in two aspects.
First, cuRobo tends to produce straight trajectories, where the hand and the in-hand object move too close to obstacles (see Fig. 9(b,c-2)). Also, as cuRobo is designed for rigid-body grippers, the object tends to produce minimal rotation after grasping, again resulting in little space between the object and the environment. In contrast, our planner, learned from kinematic features, employs dexterous strategies to produce human-like trajectories; for instance, when grasping a spatula out from a drawer, it makes use of the finger degrees of freedom for rotating the in-hand object while pulling it out vertically, thereby reducing the swept volume (see Fig. 9 (a-1)).
Secondly, cuRobo can only produce hand trajectories with a fixed base, resulting in a lack of coordination of the locomotion with the pick-and-place task; including the base translation and orientation with the minimum jerk criteria would be difficult, as human models require stepping, causing many local minima. Indeed, producing human-like motion with optimization in canonical space requires a significant amount of cost/reward tuning in high dimensions, which is not practical. In contrast, our data-driven model motion can produce well-coordinated motions that are well-suited for the condition represented by the latent code $\mathbf{z}$ and the condition parameters $\mathbf{c}$, eliminating the fixed hand base to achieve body-level guidance during its trajectory prediction.
Comparison to Diffusion Policy
Although learned from the same kinematic dataset, we find that the Diffusion Policy struggles to robustly produce realistic trajectories in test scenes due to the sparse dataset with high diversity. It suffers from over-straight trajectories (see Fig. 9(a,b-3) blue curves), irregular wrist orientation (see Fig. 9(b,c-3) and sudden change in height that result in heavy placement (see  Fig. 9(d-3)).
Table 4. Ablation study and comparison to the human motion data. The unsmoothness is computed by averaging the accelerations among all joints in $cm/s^{2}$. The sliding sums up the two foot velocities during ground contact by $cm/s$. The rooted mean-square curvature (RMSC) of the wrist trajectory is averaged among all the approaching/leaving cases in $cm^{-1}$.
@@
Quantitative Evaluation: The quantitative results in Tab. 4 also validate that our implicit neural trajectory planner can produce body motion more similar to human movement, resulting in an overall 11% lower Fréchet distance to the dataset compared with the motion following the cuRobo trajectory. Moreover, the improved motion smoothness and reduced foot sliding illustrate that our generated trajectory can be better tracked by the character, enabling it to achieve interactions with higher body coordination, benefiting from the aforementioned strengths of our kinematic data-based planner.
Figure 10. Figures from (a) to (d) sequentially show the planning of approaching a bottle with field deformation under a standard-sized shelf, an 80% narrow (29.5cmcm in width) shelf, a deeply placed case in the narrow shelf, and a deeply placed case with another bottle in a narrow shelf.
Figure 11. Visualization of the implicit neural representation of the layouts and approaching trajectories uniformly sampled from the training dataset of Sec. 4.2. We project the latent vector $\mathbf{z}$ of each 3-channel field $\mathbf{D}$ to 2D by t-SNE, discarding the condition labels $\mathbf{c}$. The latent space automatically clusters different types of furniture, while similar layouts are continuously distributed. The projection of all the unseen layouts (cases in grey plots) is mapped to those of similar scenes (in orange).
@@ 8.2.2. Evaluation Under Scene Generalization
Although our planner was only trained on standard-sized furniture in our motion capture scenes, it learns the spatial relationships between the trajectories and the environment, and can produce trajectories in novel scenes that preserve such relationships. In Fig. 10, we show snapshots of the pick-and-place motions on shelves, along with the adapted trajectories in blue and the predicted time-of-arrival field $\bm{D}_{toa}$ depicted in green. Under a standard-sized shelf with an ordinary placement (Fig. 10(a)), there is ample space for the right hand to reach the goal from the right side. When the shelf’s size is narrowed to 80% of the training size (see Fig. 10(b)), the character instead reaches from the front and passes through the small space on the right to grasp the bottle. Further, after setting the bottle’s position deeper into the narrow shelf (Fig. 10(c)), the character moves its wrist further to the left before reaching for the object. Upon adding another bottle (Fig. 10(d)), the system further adjusts the hand trajectory to avoid any collision. These experiments demonstrate that our designed three-channel fields of similar layouts can be well-mapped to a continuous latent space by the auto-decoder, enabling our planner to generalize to novel scenes.
@@
To further illustrate the generalization capability, we analyze the latent space of our auto-decoder. Similar to how DeepSDF exhibits continuous latent representations for diverse shapes, our auto-decoder maps diverse scene–trajectory pairs to the latent space. We project the latent codes onto a plane using t-SNE (Cieslak et al., 2020) and visualize them in Fig. 11. The projected scenes are well clustered, with unseen test scenes mapped adjacent to similar ones. This distribution reveals that even when training on scenes with various heterogeneous furniture together in one model, the auto-decoder structure effectively forms a continuous and smooth latent space for all types of scenes. This allows our runtime optimization to stably find the code that best suits the test environment and conditions, mapping it to a 6D trajectory within the distribution of plausible manipulations.
@@
To quantitatively evaluate the generalization capability of different trajectory planners, we prepare an extended test dataset composed of our test data and those from indoor scene databases  [^2] https://assetstor.unity.com/packages/3d/props/interior/kitchen-props-pack-pbr-204679; featuring various obstacle shapes, layout differences, and placement variations. Similar to Fig. 10, we synthesize unseen layouts by randomly shifting each obstacle object on/inside the involved furniture within 5cmcm, rotating them within $\pm 90^{\circ}$, and re-scaling them within $0.8\sim 1.25\times$. Also, we reshape the containers to new widths, depths, and heights ranging from 70% to 150% of their original dimensions, and remain the pickable cases.
We compare the trajectory un-smoothness (measured by the mean acceleration), the remaining safety distance to obstacles from the hand and object, and the success rate among the results from our planner, the NVIDIA cuRobo planner, and the Diffusion Policy, as the examples presented in Tab. 5. Our evaluation benchmark includes 168 densely packed layouts, which will be released alongside our Unity3D project.
As shown in Tab. 5 part 2, our SE(3) trajectory planner achieves a high success rate over 96%, with a 4.5% higher safety distance than the cuRobo planner and 28.4% higher than the Diffusion Policy, along with significantly better motion smoothness. These results validate the benefits of our planner’s strong generalization capability and more human-like pick-and-place trajectories. The success rate of cuRobo also decreases when the scene is highly cluttered and requires more adjustments to navigate through narrow regions. In these cases, employing a highly precise geometric approximation in cuRobo fails to converge to a feasible solution, while using a coarser geometric approximation relaxes constraints excessively, leading to collision-prone policies.
Qualitative Evaluation
@@ 8.2.3. Ablation Study on Distance Field Channels
We evaluate the effectiveness of our three-field setting by comparing it with a simplified two-channel auto-decoder. In these alternative models, we only keep the goal-reaching distance $D_{t}$ (as INTP-2Ch-$D_{t}$) or we merge the $D_{t}$ together with the obstacle distance $D_{o}$ into one channel $D_{o^{\prime}}$ (as INTP-2Ch-$D_{o^{\prime}}$) by the unsigned/signed distance to the nearest obstacle/target object surface, respectively. As shown in Tab. 5, the performance decreases significantly when using only two channels. This result confirms that having separate and complete channels for $D_{t}$ and $D_{o}$ is beneficial for high-quality trajectory predictions. Specifically, the $D_{t}$ field away from the object helps to direct the hand to approach, while the $D_{o}$ field contributes to collision avoidance.
@@
Figure 12. Ablation study for our DeepPhase interaction controller in typical full-body interactions. We visualize the hand trace for the target approaching at (a,b), where the trajectory pieces with high curvature are highlighted by red heatmaps. In line (c), we show the bipedal trace of a pivot turning case, and line (d) shows a side-stepping case, where the skeletons from green to blue are from sequentially generated key-frames where both feet are on the ground, with red curves highlighting the sliding pieces. We annotate each approaching trajectory with its RMS curvature in $cm^{-1}$, the mean sliding during stepping in $cm/s$, and the duration time in seconds.
@@ 8.3. Evaluation of the Pick-and-Place Motions
8.3.1. Ablation Study of the Deep Phase Interaction Controller
We evaluate the effectiveness of our goal-driven character controller by examining 3 variants of our DeepPhase interaction controller, including a model that drives the character solely based on the goal key joints (w/o any goal phase), a model that only uses matched goal phase features obtained from the dataset (w/o predicted goal phase), and a model where the network is only fed with goal phase features auto-regressively predicted by itself (w/o matched goal phase). In addition to testing in the general scene described earlier (see Footnote 2), we use another test scene from our dataset composed of 20 lightweight objects and 9 pieces of furniture. A long motion sequence over 3 minutes is constructed by the user through sequential clicks on objects and furniture. We compute the smoothness, foot skating artifacts, and the Fréchet distance to the dataset (see Tab. 4). The results show that the use of the goal phase improves the accuracy of all measures.
@@ 8.3.1. Ablation Study of the Deep Phase Interaction Controller
When no goal phase feature is given, the legs fail to produce effective stepping to let the key joints track their goals, resulting in larger foot skating (see Tab. 4). This is evident in complex locomotion like pivot turning (see Fig. 12(c-4)). Also, the upper body can struggle to recover from challenging previous poses without a goal cue, as shown in Fig. 12(d-4), where the uncoordinated body causes increased foot sliding.
When the matched goal phase is used instead of that predicted by the network, the system suffers from poor guidance to the goal. As the matched goal phase corresponds to sparse keyframes, it cannot guide the character to conduct efficient stepping for reaching a state that is spatio-temporally apart (see Fig. 12(c-3)). Also, the deviation of the current state and the matched state leads to biased directional guidance, requiring the character to take extra steps/time for reaching the goal (see Fig. 12(d-3)).
@@
Furthermore, the trajectory of the end-effector exhibits higher motion realism and smoothness, accompanied by body coordination when using the filtered goal phase. Without the goal cue in the frequency domain, tracking only the three key joint goals spatially leads the model to focus excessively on tracking the hip and the non-manipulating hand goals, especially when the manipulation takes place away from the torso. This causes large deviations when approaching the hand goals (see Fig. 12(b-4)) and results in an unsmooth trajectory of the manipulating hand.
Even provided with the matched goal phase priors, the system can produce unsmooth wrist motions when the matched prior suddenly switches from left-hand grasping to right-hand grasping—for example, when the main hand starts to reach out for the target object after the auxiliary hand finishes opening the cabinet, as shown in Fig. 12(b-3), and significant direction adjustments can also occur (see Fig. 12(a-3)).
While the approaching motion toward the goal phase estimated directly by the MoE network slightly improves curvature results (see Fig. 12(a-2, b-2)), it still suffers from discrete switches of the goal phase. When moving the hand goal spatially along the planned approaching trajectory, the predicted goal phase updates rapidly to catch up with the goal. In case (b-2), the arm immediately stretches out to reach the target object even before the character is close enough, producing an awkward motion due to overstretching.
@@
Figure 13. Limited performance on generating coordinate full-body collision-free interaction under novel scenes.